OpenStudy (anonymous):

The distance between the points (3,3) and (2,1) is: a) 29 b) 5 c) sqrt of 29 d) sqrt of 5 e) 1

OpenStudy (r):

\[D(P1,P2)=\sqrt{(3-2)^{2}+(3-1)^{2}}\]

OpenStudy (anonymous):

so the \[\sqrt{5} ?\]

OpenStudy (r):

\[\sqrt{(1)^{2}+(2^{2})}=\sqrt{5}\]

OpenStudy (anonymous):

thank you

OpenStudy (r):

Yes you got the correct answer.

OpenStudy (anonymous):

The general equation to calculate a distance between two points, say A and B is :\[AB = \sqrt{(x _{b}-x_{a})^{2}+(y _{b}-y_{a})^{2}}\] This also applies to points in space, just be sure to include the z component in the sum.